OpenStudy (anonymous):

For which system of equations is (5, 3) the solution? A. 3x – 2y = 9 3x + 2y = 14 B. x – y = –2 4x – 3y = 11 C. –2x – y = –13 x + 2y = –11 D. 2x – y = 7 2x + 7y = 31

OpenStudy (solomonzelman):

5 is the x value and 3 is the y value plug in 5 for x and 3 for y to see which of the options works.

OpenStudy (anonymous):

its not b

OpenStudy (solomonzelman):

true it is not b.

OpenStudy (anonymous):

I don't know this one

OpenStudy (solomonzelman):

A. 3x – 2y = 9 3x + 2y = 14 gives 3(5) – 2(3) = 9 ---> 15-6=9 ✓ 3(5) + 2(3) = 14 ---> 15+6=14 X C. –2x – y = –13 x + 2y = –11 gives -2(5) – (3) = 9 ---> -10-3=-13 ✓ (5) + 2(3) = -11 ---> 5+6=-11 X D. 2x – y = 7 2x + 7y = 31 gives 2(5) –(3) = 7 ---> 10-3=7 ✓ 2(5) + 7(3) = 31 ---> 10+21=31 ✓

OpenStudy (anonymous):

thank you so much fan and metal can u help with more?

OpenStudy (solomonzelman):

Please ask one question per thread. I am not concerned about medals, but please...

OpenStudy (solomonzelman):

close this one and ask a new one.

OpenStudy (solomonzelman):

Also, for a future reference: you don't need to read, because all of this will come to you anyway (after a certain time). you can find any section by: 1) clicking "more subjects" (the top left corner, towards the center) 2) typing in the name of the section (name of a section can be "history" Mathematics and other section for each unique area of studying). You can FAN someone by: 1) clicking on the user's name 2) clicking "fan" on the bottom right corner of user's box. ( you can also write a testimonial by clicking "add testimonial" and typing something in testimonial box) you can mention someone in your question (or in any other question) by writing @ in front of their name, just like I would mention you... LOOK ! .... @math202 (they would get a notification that a user mentioned them in a reply) you can message someone by: 1) clicking on the someone's username 2) clicking "Send Message" 3) type your message press submit, and wait till the messages submits. If you have any questions you can message me or anyone else on this website. We will be happy to respond !